# Council Mode: Mitigating Hallucination and Bias in LLMs via Multi-Agent Consensus

### 03 Apr 2026 v1

@@ 1 Introduction
In this paper, we introduce the Council Mode, a structured multi-agent consensus framework designed to mitigate hallucinations and biases in LLM outputs. Rather than relying on a single model—whether dense or MoE-based—the Council Mode queries $N$ architecturally diverse frontier models in parallel and employs a dedicated synthesis model to aggregate their responses through a formalized consensus protocol. The synthesis process explicitly categorizes the combined knowledge into four structured sections: consensus points, disagreements, unique findings, and a comprehensive analysis. This structured approach ensures that the final output is not merely a majority vote but a nuanced integration that preserves valuable minority insights while filtering out individual model hallucinations.
@@ 3.2 Phase 2: Parallel Expert Generation
For non-trivial queries ($T(q)=1$), the system dispatches the prompt to $N$ distinct expert models simultaneously. Let $\mathcal{E}=\{E_{1},E_{2},\dots,E_{N}\}$ be the set of expert models. Each expert independently generates a response:
@@
where $\mathcal{H}$ is the conversation history and $\mathcal{W}_{i}$ represents the optional web search results retrieved by expert $i$. The parallel dispatch is implemented via asynchronous concurrent execution (Promise.all() in our JavaScript implementation), ensuring that the total latency is bounded by the slowest expert rather than the sum of all expert latencies:
@@ 3.3 Phase 3: Consensus Synthesis
The consensus section identifies factual claims $f$ that are supported by all $N$ experts:
@@ 3.4 Theoretical Analysis: Why Council Reduces Hallucination
We provide a probabilistic argument for why the Council Mode reduces hallucination. Let $p_{i}$ denote the probability that expert $E_{i}$ hallucinates on a given claim. Assuming independence between experts (justified by their architectural diversity), the probability that all $N$ experts hallucinate on the same claim is:

### 21 Apr 2026 v2

@@ 1 Introduction
In this paper, we introduce the Council Mode, a structured multi-agent consensus framework designed to mitigate hallucinations and biases in LLM outputs. Rather than relying on a single model—whether dense or MoE-based—the Council Mode queries NN architecturally diverse frontier models in parallel and employs a dedicated synthesis model to aggregate their responses through a formalized consensus protocol. The synthesis process explicitly categorizes the combined knowledge into four structured sections: consensus points, disagreements, unique findings, and a comprehensive analysis. This structured approach ensures that the final output is not merely a majority vote but a nuanced integration that preserves valuable minority insights while filtering out individual model hallucinations.
@@ 3.2 Phase 2: Parallel Expert Generation
For non-trivial queries ($T(q)=1$), the system dispatches the prompt to NN distinct expert models simultaneously. Let $\mathcal{E}=\{E_{1},E_{2},\dots,E_{N}\}$ be the set of expert models. Each expert independently generates a response:
@@
where $\mathcal{H}$ is the conversation history and $\mathcal{W}_{i}$ represents the optional web search results retrieved by expert ii. The parallel dispatch is implemented via asynchronous concurrent execution (Promise.all() in our JavaScript implementation), ensuring that the total latency is bounded by the slowest expert rather than the sum of all expert latencies:
@@ 3.3 Phase 3: Consensus Synthesis
The consensus section identifies factual claims $f$ that are supported by all NN experts:
@@ 3.4 Theoretical Analysis: Why Council Reduces Hallucination
We provide a probabilistic argument for why the Council Mode reduces hallucination. Let $p_{i}$ denote the probability that expert $E_{i}$ hallucinates on a given claim. Assuming independence between experts (justified by their architectural diversity), the probability that all NN experts hallucinate on the same claim is: